A golf ball with mass 45.90 g and diameter 42.60 mm is struck such that it moves with a speed of 51.85 ⁄ and rotates with a frequency of 2857rpm. What is the kinetic energy of the golf ball?


1
Expert's answer
2021-03-31T16:14:13-0400

The kinetic energy of the golf ball can be found as follows:


"KE=KE_{translational}+KE_{rotational}."

We can find the translational kinetic energy of the golf ball as follows:


"KE_{translational}=\dfrac{1}{2}mv^2,""KE_{translational}=\dfrac{1}{2}\cdot0.0459\ kg\cdot(51.85\ \dfrac{m}{s})^2=61.7\ J."

We can find the rotational kinetic energy as follows:


"KE_{rotational}=\dfrac{1}{2}I\omega^2=\dfrac{1}{2}\cdot\dfrac{2}{5}\cdot mr^2\omega^2,"

"KE_{rotational}=\dfrac{1}{5}\cdot0.0459\ kg\cdot(0.0213\ m)^2\cdot(2857\ \dfrac{rev}{min}\cdot\dfrac{1\ min}{60\ s}\cdot2\pi\ rad)^2,"


"KE_{rotational}=0.372\ J."

Then, we can calculate the kinetic energy of the golf ball:


"KE=KE_{translational}+KE_{rotational}=61.7\ J+0.372\ J=62.1\ J."